Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
Tringali, Fiori, De Rossi, Gosselin - 17:05 Tuesday 12 March 2024 (63577) Print this report
Shaker installation on laser bench

During the maintenance, we installed a shaker on the laser bench, see the attached Figures 1 and 2.
We checked the optimal parameter configuration to have the maximum displacement of the bench:

- freq. line injected at 17.9 Hz (laser bench horizontal resonance)
- voltage amplitude 0.1 V.

Looking at the episensor ENV_LB_SEIS_*, we achieved a spectral displacement of the laser bench at 17.9 Hz:

channel direction m s^{-2}/\sqrt{Hz} \mu m/\sqrt{Hz}
x  ~1e-2 ~0.8 
y ~1e-3 ~0.08
z ~4e-2 ~3.2

We checked that this injected line was not visible on EIB, Figure 3. 
We also tried to inject lines in the freq. range (1-5) Hz but the maximum displacement achieved was lower by a factor of ten.

Images attached to this report
Comments to this report:
derossi, spinicelli, tringali - 13:10 Thursday 04 April 2024 (63838) Print this report

On the March 12th, a shaker has been installed on the Laser Bench to perform sismic noise injection. After those studies done on the 14th, the electronics that controls the shaker (deployed on the ground) has been left ON (with 0V of input voltage) to have the possibility to further perform other remote injections. 
However, entering in the laser lab few days later, we noticed a slightly higher acoustic noise than usual. We decided then to switch OFF that electronics, and indeed the noise reduced.

Looking at the environmental sensors present in laser lab, we found this noise visible as both high frequency broad noise (>200Hz)  and with some lines a low frequency (~45Hz and harmonics).
Moreover, we see it on acoustic (see fig. 1-3, 5) and sismic (fig. 4) sensors. While the noise is visible on both benches, the effect (seismic and acoustic) on EIB is really faint (see fig 4-5).

Nevertheless, we didn't find any noticeble effect on Hrec.  

Images attached to this comment
Tringali, Melo - 11:51 Tuesday 11 June 2024 (64480) Print this report

This morning, we removed the shaker from the laser bench.

Images attached to this comment
Search Help